# Improving Low-Resource Sequence Labeling with Knowledge Fusion and Contextual Label Explanations

### v2

@@ Abstract
Sequence labeling remains a significant challenge in low-resource, domain-specific scenarios, particularly for character-dense languages. Existing methods primarily focus on enhancing model comprehension and improving data diversity to boost performance. However, these approaches still struggle with inadequate model applicability and semantic distribution biases in domain-specific contexts. To overcome these limitations, we propose a novel framework that combines an LLM-based knowledge enhancement workflow with a span-based Knowledge Fusion for Rich and Efficient Extraction (KnowFREE) model. Our workflow employs explanation prompts to generate precise contextual interpretations of target entities, effectively mitigating semantic biases and enriching the model’s contextual understanding. The KnowFREE model further integrates extension label features, enabling efficient nested entity extraction without relying on external knowledge during inference. Experiments on multiple Chinese domain-specific sequence labeling datasets demonstrate that our approach achieves state-of-the-art performance, effectively addressing the challenges posed by low-resource settings.
@@
Peichao Lai1, Jiaxin Gan2, Feiyang Ye2, Wentao Zhang1, Fangcheng Fu1, Yilei Wang2, Bin Cui1,* 1School of Computer Science, Peking University, 2College of Computer and Data Science, Fuzhou University *bin.cui@pku.edu.cn
@@ 1 Introduction
Sequence labeling is a fine-grained information extraction (IE) task that includes sub-tasks such as named entity recognition (NER) and part-of-speech (POS) tagging, playing a critical role in various downstream natural language processing (NLP) applications. However, in low-resource scenarios, sequence labeling for character-dense languages like Chinese presents unique challenges. The absence of explicit word boundaries makes accurate label inference more complex, while the scarcity of domain-specific data further limits the model’s ability to effectively capture label distributions.
Previous studies on enhancing sequence labeling in low-resource with character-dense languages can be broadly classified into two streams: (1) Model-Centric Optimization. These methods focus on enhancing model’s comprehension to detect implicit word boundaries and contextual signals through feature engineering. For instance, lexical features are injected via lexicon matching networks (Zhang and Yang, 2018a; Li et al., 2020; Liu et al., 2021; Wu et al., 2021) or prompt templates (Ma et al., 2022b; Shen et al., 2023; Chen et al., 2021b) to strengthen entity boundary detection, while others employ knowledge transfer mechanisms like Gaussian embeddings (Si et al., 2024; Das et al., 2022), prompt-based metrics (Chen et al., 2023; Lai et al., 2022) and contrastive learning (Huang et al., 2022; Zhang et al., 2024) to distill knowledge into target domains. (2) Data-Centric Augmentation. Meanwhile, other studies concentrate on using data augmentation through altering entity label information (Hu et al., 2023; Yang et al., 2018) and extracting knowledge from the external environment (Cai et al., 2023; Chen et al., 2021a; Yaseen and Langer, 2021) to enrich the dataset. With the advent of large language models (LLMs), recent findings leverage their generative capabilities to enhance the diversity of entity and sentence synthesis (Kang et al., 2024; Ye et al., 2024).
However, critical limitations persist when applying these solutions to specialized domains: (1) Limited Model Applicability. Existing model-centric approaches for character-dense languages often struggle to integrate diverse feature types and labels effectively, restricting the flexibility and richness of feature injection. Moreover, they face challenges in handling nested entities, further constraining their adaptability. Additionally, most approaches depend on rigid feature injection frameworks and complex input configurations to enhance word boundary detection, leading to increased reliance on supplementary structures during inference and higher usage costs. (2) Variability in Label Distribution. Existing data-centric augmentation techniques suffer from domain distribution bias. Differences in entity type definitions and sample semantics across domains lead to misaligned label priors and entity contexts, significantly compromising both the quality of augmented data and zero-shot generalization. As illustrated in Figure 1, the aforementioned challenges of structural inflexibility and distribution discrepancy collectively constrain current methods’ practical efficacy, motivating our investigation into a unified framework that simultaneously resolves model architecture limitations and domain adaptation barriers.
@@
Figure 1: Distinctions between our method and existing methods in terms of model and data.
@@
In this task, we adopt two key strategies for improving low-resource sequence labeling in character-dense languages: (i) enhancing the utilization of non-entity features and (ii) improving the model’s contextual understanding of target entities. To achieve these objectives, we propose a novel LLM-based data augmentation framework. Our approach begins by designing extraction prompts to identify and extract potential non-target entity features from the input text, thereby maximizing the utilization of non-entity features. To address the issue of limited model applicability, we introduce a span-based model, Knowledge Fusion for Rich and Efficient Extraction (KnowFREE) model, which supports nested label annotations and integrates extension label features by a local multi-head attention module. Unlike prior methods, KnowFREE learns rich contextual information during training without relying on external knowledge at inference. To tackle the issue of variability in label distribution, we incorporate explanation prompts inspired by label explanation techniques (Golde et al., 2024; Yang and Katiyar, 2020; Ma et al., 2022a), generating precise explanations for target entities in their contexts. This step further strengthens the model’s contextual understanding of target entities. This novel use of LLMs for label interpretation synthesis effectively mitigates semantic distribution bias, outperforming traditional data augmentation techniques in low-resource scenarios. We evaluate our framework on multiple Chinese domain-specific sequence labeling datasets, and experimental results demonstrate its effectiveness in addressing the challenges of low-resource settings.
@@ 2 Related Work
Span-based Sequence Labeling Methods Span-based sequence labeling methods have gained prominence for their ability to address overlapping and nested entities effectively. Early works, such as Dozat and Manning (2017); Yu et al. (2020) introduced biaffine models to capture sentence-wide structures and score span boundaries for accurate information extraction. Based on this, Su et al. (2022) proposed the Global Pointer model, optimizing the biaffine transformation’s weight matrix and bias terms to boost efficiency and precision in span-based NER. In parallel, Shen et al. (2022) introduced a parallel instance query network for simultaneous entity extraction. Then, (Yan et al., 2023) proposed a multi-head biaffine mechanism combined with CNNs to capture local span features, achieving improved performance, while (Li et al., 2022) combined bilinear classifiers with dilated convolutions post-CLN to refine span-level relation classification. For few-shot NER, (Wang et al., 2022) introduced SpanProto, which integrates prototype-based classification with a contrastive loss to effectively separate non-target spans from prototype clusters, demonstrating strong performance in low-resource scenarios.
Sequence Labeling via LLMs The advent of LLMs (OpenAI, 2023; Bai et al., 2023; Touvron et al., 2023) has introduced innovative strategies for sequence labeling. Generative approaches utilizing in-context learning (ICL) allow LLMs to perform sequence labeling tasks directly (Jiang et al., 2024). Leveraging their extensive knowledge and flexible template structures, LLMs have also shown promise in supporting cross-domain entity recognition (Nandi and Agrawal, 2024) and tackling nested entity challenges (Kim et al., 2024). In few-shot scenarios, methods like ProgGen (Heng et al., 2024) exploit LLMs to iteratively synthesize attribute-rich training data by incorporating domain-specific features. Chen et al. (2024) integrates a small span detection model with LLM validation to enhance span selection and address domain-specific challenges.
@@ 3 Method
In this section, we will introduce the knowledge enhancement workflow and the specific structure of our KnowFree model for information extraction. The overall framework is shown in Figure 2.
Figure 2: The workflow of knowledge enhancement.
@@ 3.1 Workflow of Knowledge Enhancement
Label Extension Annotation. In low-resource scenarios, non-entity segments in the sentence may contain additional non-target entity features, while data samples with flat-only entities may potentially contain nested entity information around the target entities. Leveraging these potential feature information can enhance the model’s capacity to comprehend the fine-grained semantics and the ability to distinguish entity boundaries in character-dense languages. To achieve this, we utilize the LLM’s general knowledge to generate extension entity tags, word segmentation tags, and part-of-speech (POS) tags for the source samples.
@@
We assume that accurate and diverse extension tags can significantly improve the model’s comprehension of context and its capacity for entity detection. Then, there are several issues that need to be solved in the results of extraction. (1) The uncertainty generation by LLMs leads to the presence of tags in the extension entity set in various textual formats yet denotes the same label category. Directly introducing these labels will significantly disrupt the model’s assessment of the entity type. (2) Word segmentation usually produces boundary labels with low information entropy. However, POS tagging based on LLMs sometimes has missing annotations. Notably, POS tagging labels inherently include implicit word boundary information. Therefore, combining the outputs of these two processes is expected to yield better results.
@@
where $\tilde{T}_{i}$ is the standard label, $T_{ij}$ is the synonymous label, and $m$, $r$ is the number of standard label and its corresponding synonymous labels, respectively. This method is determined by the LLM’s literal interpretation of the label, which may not accurately align the semantic spatial distribution of entities in the target domain, and therefore is not completely reliable. We further compute the vector representation of each entity-label pair by using a sentence embedding model $\mathcal{M}$, and vector set of $T_{k}$ can be represented as:
@@
$\mathbf{v}_{i}=\mathcal{M}(x_{i}),$ (5)
where the input $x_{i}$ is the concatenation of $e_{i}$ and $t_{i}$ with the template of “[$e_{i}$] is [$t_{i}$]”. The center point and mean radius of the vector set for $T_{k}$ can be calculated as:
@@
where $p$ represents Top-$p$ samples with the largest distance from $\mathbf{c}_{k}$. Then, for each standard label $\tilde{T}_{i}$, we identify the synonymous label vector set $\mathbf{V}_{i,max}$ that contains the largest number of samples and has the maximum radius, designating it as the reference vector set. Subsequently, we evaluate whether other synonymous label vector sets $\mathbf{V}_{i,j}$ satisfy the following condition:
$\|\mathbf{c}_{j}-\mathbf{c}_{max}\|\leq\epsilon\cdot r_{max}.$ (8)
If the condition holds, $T_{j}$ is merged into $\tilde{T}_{i}$; Otherwise, $T_{j}$ is treated as an independent standard label. To address the second issue, we initially compute the word segmentation set $\hat{E}^{\mathrm{seg}}$ utilizing $\mathcal{P}_{seg}$, then we employ $\hat{E}^{\mathrm{seg}}$, $\mathcal{P}_{pos}$ and the source sample as input for the LLM to generate the part of speech without omission. This effectively ensures that all words in the sample are tagged with part of speech, and also improves the diversity of word segmentation features.
Finally, we merge all the extension entities with standardized labels into the original data to obtain the fusion samples. We use our KnowFree model to first train an annotation model on the fusion samples, which can be used for entity annotation in the next pipeline.
@@
Next, we generate enriched explanations using the explanation prompts and proceed with annotating the entities in these synthetic texts through a two-branch pipeline: one for target entity annotation and the other for extension entity extraction. The frozen KnowFree model, trained in the previous pipeline, is used to annotate the target entities. For extracting extension entities, we apply the same method as in the previous pipeline. Finally, all entities are integrated into the enriched explanations to produce synthetic samples. As shown in Figure 2, we can combine fusion samples and synthetic samples and retrain on the KnowFree model to enhance its performance in the few-shot scenario.
@@ 3.2 Structure of KnowFree Model
Figure 3: The structure of KnowFree model.
To support the fusion of multi-label knowledge, our proposed model is implemented based on the biaffine structure. The structure of our KnowFree model is illustrated in Figure 3. Compared to previous approaches that involved external feature injection, our approach avoids the creation of additional feature injection modules at the input stage. Both extension entity labels and target entity labels are input together into the model for training, with their weights adjusted through hyperparameters during the loss function computation. During inference, we mask the weights of the extension entity labels, ensuring that the model only predicts target entities. This approach significantly simplifies the model while improving its overall efficiency. We additionally incorporate a local multi-head attention layer into the model to improve the interaction of multi-label features among neighborhoods during the training process.
Formally, let denote $X,\hat{X}$ as inputs of the fusion sample and the synthetic sample, respectively. The target entity spans of each sample is represented as $[s_{i},e_{i},l_{i}]$, where $s_{i},e_{i}$ are the start and end indices of the entity span, and $l_{i}$ is the label type. The extension entity span is represented as $[s_{j},e_{j},\tilde{l}_{j}]$, where $\tilde{l}_{j}$ is the extension label type. We adopt a pre-trained encoder to compute the hidden states $H\in\mathbb{R}^{L\times D}$ for each input, where $L$ is the length of the input sequence, $D$ is the dimension of hidden states. We then compute the encoding of the entity’s start and end positions:
@@
where $\mathbf{W}_{s},\mathbf{W}_{e}\in\mathbb{R}^{D\times D^{\prime}}$ are learnable weight matrics, $D^{\prime}$ is the feature hidden size, and $\sigma$ is the activation function. The biaffine matrix of spans is computed by a multi-head Biaffine decoder:
@@
where $\mathcal{F}_{\text{MHB}}$ represents the multi-head Biaffine decoder, and $H_{\text{B}}\in\mathbb{R}^{L\times L\times\tilde{D}}$, $\tilde{D}$ is the hidden size of the biaffine matrix.
To improve the interactivity between multi-label features and span neighborhoods, we introduce a local multi-head attention layer to generate a mask for local multi-head attention, ensuring that each token can only interact with its neighboring tokens:
@@
where $M\in\mathbb{R}^{L\times L}$, $\omega$ is the window size of local multi-head attention. The input feature $H_{\text{B}}$ is divided into $K$ attention heads, each with a reduced feature dimension of $\tilde{D}/K$. For each attention head, we compute the query $Q$, key $K$, and value $V$ matrices via linear projections of the input:
$Q^{(k)}=H_{\text{B}}\mathbf{W}_{Q}^{(k)},K^{(k)}=H_{\text{B}}\mathbf{W}_{K}^{(k)},V^{(k)}=H_{\text{B}}\mathbf{W}_{V}^{(k)},$
where $\mathbf{W}_{Q}^{(k)},\mathbf{W}_{K}^{(k)},\mathbf{W}_{V}^{(k)}\in\mathbb{R}^{\tilde{D}\times(\tilde{D}/{K})}$ are learnable projection matrices for the $k$-th head. We then compute the scaled dot-product attention for each attention head:
$\mathcal{A}(Q^{(k)},K^{(k)},V^{(k)},M)=\text{Softmax}\left(\frac{Q^{(k)}(K^{(k)})^{T}+M}{\sqrt{d_{k}}}\right)V^{(k)},$
where $d_{k}=\tilde{D}/{K}$ is the dimension size of the $k$-th head. The outputs from all attention heads are concatenated as:
$\mathcal{\tilde{A}}=[\mathcal{A}_{1};\mathcal{A}_{2};\ldots;\mathcal{A}_{K}],$ (14)
$H_{\text{attn}}=\text{LayerNorm}(\mathcal{\tilde{A}}),$ (15)
where $\mathcal{\tilde{A}}\in\mathbb{R}^{L\times L\times\tilde{D}}$, and the output feature is finally computed with a layer normalization. To enhance training stability and retain the initial biaffine features, we apply a residual connection followed by layer normalization:
@@
We then use the fully connection layer to map the sum of $H_{\text{G}}$ and $H_{\text{B}}$ into the number of entity tags:
@@
$\mathcal{L}=-(\sum_{\begin{subarray}{c}0\leq i,\\ j<N_{\text{tgt}}\end{subarray}}y_{i,j}\log P_{i,j}+\alpha\sum_{\begin{subarray}{c}N_{\text{tgt}}\leq i,\\ j<N_{\text{ext}}\end{subarray}}y_{i,j}\log P_{i,j}),$ (18)
where $y$ refers to the ground truth labels, $\alpha$ is the weight parameter. Similarly, we control the influence of the quantity and noise in synthetic samples with another weight parameter. The final loss is a weighted sum of the original and synthetic losses:

### v4

@@ Abstract
Sequence labeling remains a significant challenge in low-resource, domain-specific scenarios, particularly for character-dense languages. Existing methods primarily focus on enhancing model comprehension and improving data diversity to boost performance. However, these approaches still struggle with inadequate model applicability and semantic distribution biases in domain-specific contexts. To overcome these limitations, we propose a novel framework that combines an LLM-based knowledge enhancement workflow with a span-based Knowledge Fusion for Rich and Efficient Extraction (KnowFREE) model [^1] Our code: https://github.com/aleversn/KnowFREE. Our workflow employs explanation prompts to generate precise contextual interpretations of target entities, effectively mitigating semantic biases and enriching the model’s contextual understanding. The KnowFREE model further integrates extension label features, enabling efficient nested entity extraction without relying on external knowledge during inference. Experiments on multiple domain-specific sequence labeling datasets demonstrate that our approach achieves state-of-the-art performance, effectively addressing the challenges posed by low-resource settings.
@@
Peichao Lai1, Jiaxin Gan3, Feiyang Ye3, Wentao Zhang1, Fangcheng Fu1, Yilei Wang3, †, Bin Cui1,2,* 1School of Computer Science, Peking University, 2Institute of Computational Social Science, Peking University (Qingdao), 3College of Computer and Data Science, Fuzhou University *bin.cui@pku.edu.cn yilei@fzu.edu.cn
@@ 1 Introduction
Sequence labeling is a fine-grained information extraction (IE) task that includes sub-tasks such as named entity recognition (NER), word segmentation, and part-of-speech (POS) tagging, playing a critical role in various downstream natural language processing (NLP) applications.
In low-resource scenarios, sequence labeling remains a persistent challenge, primarily due to the scarcity of domain-specific data, which limits the model’s capacity to learn accurate label distributions. Moreover, character-dense languages such as Chinese pose additional difficulties, as the absence of explicit word boundaries greatly complicates label inference.
@@
Figure 1: Distinctions between our method and existing methods in terms of model-centric and data-centric.
Previous studies predominantly focus on two main directions to enhance sequence labeling in low-resource scenarios: (1) Model-Centric Optimization. These methods focus on enhancing the model’s comprehension to detect implicit word boundaries and contextual signals through feature engineering. For instance, lexical features are injected via lexicon matching networks (Zhang and Yang, 2018a; Li et al., 2020; Liu et al., 2021; Wu et al., 2021) or prompt templates (Ma et al., 2022b; Shen et al., 2023; Chen et al., 2021b; Das et al., 2023) to strengthen entity boundary or type detection. Other methods employ knowledge transfer techniques such as Gaussian embeddings (Si et al., 2024; Das et al., 2022), prompt-based metrics (Chen et al., 2023; Lai et al., 2022), and contrastive learning (Huang et al., 2022; Zhang et al., 2024) to distill knowledge into target domains. (2) Data-Centric Augmentation. Meanwhile, data-centric methods concentrate on using data augmentation through altering entity label information (Hu et al., 2023; Yang et al., 2018), back translation (Paolini et al., 2021; Yaseen and Langer, 2021), and extracting knowledge from the external environment (Cai et al., 2023; Chen et al., 2021a; Yaseen and Langer, 2021) to enrich the dataset. With the advent of large language models (LLMs), recent findings leverage their generative capabilities to enhance the diversity of entity and sentence synthesis (Kang et al., 2024; Ye et al., 2024).
However, as illustrated in Figure 1, significant limitations remain when applying these solutions to specialized domains: (1) Limited Model Applicability. Existing model-centric approaches for character-dense languages often struggle to effectively incorporate diverse feature types and label structures, limiting the flexibility and expressiveness of feature injection. These methods also face difficulties in handling nested entities, further reducing their adaptability. Moreover, many approaches rely on rigid feature integration pipelines and complex input configurations to improve word features, leading to increased reliance on supplementary structures during inference and raising deployment costs. (2) Variability in Label Distribution. Existing data-centric augmentation methods frequently suffer from domain distribution biases. Inconsistencies in entity type definitions and semantic contexts across domains lead to mismatches in label priors and entity representations, undermining the quality of synthesized data and weakening zero-shot generalization.
These challenges, including structural rigidity and distributional mismatch, collectively hinder the practical effectiveness of current methods. This motivates our development of a unified framework that addresses both architectural constraints and domain adaptation challenges in a holistic manner. In this task, we adopt two key strategies for improving low-resource sequence labeling in character-dense languages: (i) enhancing the utilization of non-entity features through the span-based model and (ii) improving the model’s contextual understanding of target entities.
@@
To achieve these objectives, we propose a novel LLM-based data augmentation framework. Our approach begins by designing extraction prompts to identify and extract informative non-target entity features from the input text, thereby maximizing the utilization of non-entity information. To address the issue of limited model applicability, we introduce a span-based model called Knowledge Fusion for Rich and Efficient Extraction (KnowFREE), which supports nested entity annotation and integrates extension label features through a local multi-head attention module. Unlike previous methods, KnowFREE captures rich contextual representations during training without relying on external knowledge at inference time. To tackle the issue of variability in label distribution, we incorporate explanation prompts inspired by label explanation techniques (Golde et al., 2024; Yang and Katiyar, 2020; Ma et al., 2022a), enabling the generation of precise, context-aware explanations for target entities. This enhances the model’s contextual understanding and mitigates semantic distribution biases. By leveraging LLMs for label interpretation synthesis, our framework outperforms other related data augmentation techniques in low-resource settings. We evaluate it on multiple Chinese and English domain-specific sequence labeling datasets, and experimental results demonstrate its effectiveness in overcoming the key limitations of low-resource scenarios.
@@ 2 Related Work
Span-based Sequence Labeling Methods Span-based sequence labeling methods have gained prominence for their ability to address overlapping and nested entities effectively (Yang and Tu, 2022; Fu et al., 2021). Early works, such as Dozat and Manning (2017); Yu et al. (2020) introduced Biaffine models to capture sentence-wide structures and score span boundaries for accurate information extraction. Based on this, Su et al. (2022) proposed the Global Pointer model, optimizing the Biaffine transformation’s weight matrix and bias terms to boost efficiency and precision in span-based NER. In parallel, Shen et al. (2022) introduced a parallel instance query network for simultaneous entity extraction. Then, (Yan et al., 2023) proposed a multi-head Biaffine mechanism combined with CNNs to capture local span features, achieving improved performance, while (Li et al., 2022) combined bilinear classifiers with dilated convolutions post-CLN to refine span-level relation classification. For few-shot NER, (Wang et al., 2022) introduced SpanProto, which integrates prototype-based classification with a contrastive loss to effectively separate non-target spans from prototype clusters, demonstrating strong performance in low-resource scenarios.
Sequence Labeling via LLMs
Recent advances in LLMs (OpenAI, 2023; DeepSeek-AI et al., 2025; Touvron et al., 2023) have introduced new paradigms for sequence labeling. Generative methods based on in-context learning (ICL) allow LLMs to perform labeling tasks directly without task-specific fine-tuning (Jiang et al., 2024). In zero-shot settings, InstructUIE (Wang et al., 2023) adopts a single-turn instruction framework across diverse IE tasks, UniversalNER (Mayhew et al., 2024) demonstrates improved performance by querying one entity type at a time, and GoLLIE (Sainz et al., 2024) enhances generalization via structured code-style prompting. LLMs have also shown promise in handling cross-domain and nested entity recognition (Nandi and Agrawal, 2024; Kim et al., 2024). In parallel, LLM-based data augmentation strategies synthesize high-quality training data by injecting domain-specific features (Ye et al., 2024; Heng et al., 2024), while others combine lightweight span detectors with LLM validation to improve span selection in specialized domains (Chen et al., 2024).
@@ 3 Method
In this section, we will introduce the knowledge enhancement workflow in § 3.1 and the specific structure of our sequence labeling KnowFREE model in § 3.2. The overall framework is shown in Figure 2.
Figure 2: The workflow (top) and detailed pipeline structure (bottom) of our knowledge enhancement framework. Pipeline 1 generates extension entities to enhance the performance of KnowFREE, while Pipeline 2 synthesizes additional training samples and entities. We then use a frozen KnowFREE model to annotate target entities within these synthetic samples.
@@ 3.1 Workflow of Knowledge Enhancement
Label Extension Annotation. In low-resource scenarios, non-entity segments in the sentence may contain additional non-target entity features, while data samples with flat-only entities may potentially contain nested entity information around the target entities. Leveraging this potential feature information can enhance the model’s capacity to comprehend the fine-grained semantics and the ability to distinguish entity boundaries in character-dense languages. To achieve this, we utilize the LLM’s general knowledge to generate extension entity tags, word segmentation tags, and part-of-speech (POS) tags for the source samples.
@@
We assume that accurate and diverse extension tags can significantly improve the model’s comprehension of context and its capacity for entity detection. Then, there are several issues that need to be solved in the results of extraction. (1) The inherent uncertainty in LLM generation often produces extension entity tags in multiple textual forms that nonetheless correspond to the same label category, a phenomenon known as Surface Form Competition (Holtzman et al., 2021). Directly introducing these labels will significantly disrupt the model’s assessment of the entity type. (2) Word segmentation usually produces boundary labels with low information entropy. However, POS tagging based on LLMs sometimes has missing annotations. Notably, POS tagging labels inherently include implicit word boundary information. Therefore, combining the outputs of these two processes is expected to yield better results.
@@
where $\tilde{T}_{i}$ is the standard label, $T_{ij}$ is the synonymous label, and $m$, $r$ is the number of standard label and its corresponding synonymous labels, respectively. This method is determined by the LLM’s literal interpretation of the label, which may not accurately align the semantic spatial distribution of entities in the target domain, and therefore is not completely reliable. We further compute the vector representation of each entity-label pair by using a sentence embedding model $\mathcal{M}$, and the vector set of $T_{k}$ can be represented as:
@@
where $\mathbf{v}_{i}$ refers to $\mathcal{M}(x_{i})$, $x_{i}$ is the concatenation of $e_{i}$ and $t_{i}$ with the template of $e_{i}$] is $t_{i}$]”. The center point and mean radius of the vector set for $T_{k}$ can be calculated as:
@@
where $p$ denotes the Top-$p$ samples that exhibit the greatest distance from $\mathbf{c}_{k}$. For each standard label $\tilde{T}_{i}$, we identify the synonymous label vector set $\mathbf{V}_{i,max}$ that contains the largest number of samples and has the largest radius, designating it as the reference vector set. We then evaluate whether each remaining synonymous label vector set $\mathbf{V}_{i,j}$ satisfies the condition $\|\mathbf{c}_{j}-\mathbf{c}_{max}\|\leq\epsilon\cdot r_{max}$. If the condition is met, $T_{j}$ is merged into $\tilde{T}_{i}$; otherwise, $T_{j}$ is treated as an independent standard label.
To address the second issue, we first compute the word segmentation set $\hat{E}^{\mathrm{seg}}$ using $\mathcal{P}_{seg}$. Then, we combine $\hat{E}^{\mathrm{seg}}$, $\mathcal{P}_{pos}$, and the original source sample as input to the LLM to generate part-of-speech tags without omissions. This approach ensures comprehensive POS tagging for all words in the sample while enhancing the diversity of word segmentation features.
Finally, we merge all the extension entities with standardized labels into the original data to obtain the fusion samples. We use our KnowFREE model to first train an annotation model on the fusion samples, which is then employed to annotate target entities in the synthetic samples generated by the subsequent pipeline.
Figure 3: The architecture of the KnowFREE model. The span logits corresponding to the extension entity labels are ignored during inference. Matrices like (a), (b), (c), and (d) represent the span logits for each label type.
@@
Next, we generate enriched explanations using explanation prompts and annotate entities in the synthetic texts through two branches: one for target entity annotation and the other for extension entity extraction. The frozen KnowFREE model trained in Pipeline 1 is used to annotate target entities, while extension entities are extracted by reapplying Pipeline 1 to the synthetic texts. Finally, the fusion entities of synthetic samples are obtained by integrating both extension and target entities. As shown in Figure 2, we combine these fusion samples with synthetic samples and retrain the KnowFREE model to further improve its performance in low-resource scenarios.
@@ 3.2 Structure of KnowFREE Model
To support the fusion of multi-label knowledge, the KnowFREE model is built upon a Biaffine architecture, as illustrated in Figure 3. Unlike previous methods that rely on external feature injection, our approach eliminates the need for additional injection modules at the input stage.
Formally, let denote $X,\hat{X}$ as inputs of the fusion sample and the synthetic sample, respectively. The target entity spans of each sample is represented as $s_{i},e_{i},l_{i}]$, where $s_{i},e_{i}$ are the start and end indices of the entity span, and $l_{i}$ is the label type. The extension entity span is represented as $s_{j},e_{j},\tilde{l}_{j}]$, where $\tilde{l}_{j}$ is the extension label type. We adopt a pre-trained encoder to compute the hidden states $H\in\mathbb{R}^{L\times D}$ for each input, where $L$ is the length of the input sequence, $D$ is the dimension of hidden states. We then compute the encoding of the entity’s start and end positions:
@@
where $\mathbf{W}_{s},\mathbf{W}_{e}\in\mathbb{R}^{D\times D^{\prime}}$ are learnable weight matrics, $D^{\prime}$ is the feature hidden size, and $\sigma$ is the activation function. The Biaffine matrix of spans is computed by a multi-head Biaffine decoder $\mathcal{F}_{\text{MHB}}$ (Yu et al., 2020):
@@
where $H_{\text{B}}\in\mathbb{R}^{L\times L\times\tilde{D}}$, $\tilde{D}$ is the hidden size of the Biaffine matrix. To improve the interactivity between multi-label features and span neighborhoods, we introduce a local multi-head attention layer to generate a mask for local multi-head attention. Each token is restricted to attend only within a local window of size $\omega$ through a masking scheme:
@@
where $M\in\mathbb{R}^{L\times L}$, $i$, $j$, refer to the row and column indices in $M$. For input features $H_{\text{B}}$, the attention computation follows the standard multi-head paradigm with $\mathcal{K}$ heads, but incorporates the local mask $M$:
[…]
The outputs $H_{\text{attn}}$ from all heads are concatenated and processed with layer normalization. We maintain training stability through residual connections:
@@
We then use the fully connected layer to map the sum of $H_{\text{G}}$ and $H_{\text{B}}$ into the number of entity tags:
@@
[…]
where $y$ refers to the ground truth labels, $\alpha$ is the weight parameter, and $i$, $j$ denote the indices of different label types. Similarly, we control the influence of the quantity and noise in synthetic samples with another weight parameter. The final loss is a weighted sum of the original and synthetic losses:
@@
During inference, the weights associated with the extension entity labels are masked, ensuring that the model exclusively predicts the target entities. This design simplifies the overall architecture while enhancing model efficiency.